# Rethinking Positive Pairs in Contrastive Learning

### v1

@@ Appendix C Self-supervised Learning
Table 4 compares the KNN (k=10) performance across 4 classification tasks. The introduction of supervision significantly enhances performance on Cars and Food datasets, leveraging transferable information. However, it slightly diminishes performance on CIFAR10 and CIFAR100. Notably, Hydra outperforms both its unsupervised and supervised counterparts with significant improvements, particularly on the Cars dataset. These results demonstrate that common features across classes, as captured by our method, improve transferability.
Hydra with Pseudo Labels
We explore the effectiveness of Hydra in a self-supervised setting using pseudo labels. We apply K-Means clustering to generate 1,000 pseudo labels based on features extracted from SimCLR, which yields low-quality supervision. We then train Supcon and Hydra for 200 epochs using these pseudo labels. The results in Table 4 show a significant performance drop on Cars, reflecting the weak supervision provided by SimCLR. However, Hydra consistently outperforms Supcon across all four tasks, indicating that additional information is learned from arbitrary pairs even with bad pseudo labels.
@@ D.2 Training Detail
config Hydra Hydra Hydra-MoCo

### v2

@@ Appendix C Self-supervised Learning
Table 4 compares the KNN (k=10) performance across 4 classification tasks. The introduction of supervision significantly enhances performance on Cars and Food datasets, leveraging transferable information. However, it slightly diminishes performance on CIFAR10 and CIFAR100. Notably, SimLAP outperforms both its unsupervised and supervised counterparts with significant improvements, particularly on the Cars dataset. These results demonstrate that common features across classes, as captured by our method, improve transferability.
SimLAP with Pseudo Labels
We explore the effectiveness of SimLAP in a self-supervised setting using pseudo labels. We apply K-Means clustering to generate 1,000 pseudo labels based on features extracted from SimCLR, which yields low-quality supervision. We then train Supcon and SimLAP for 200 epochs using these pseudo labels. The results in Table 4 show a significant performance drop on Cars, reflecting the weak supervision provided by SimCLR. However, SimLAP consistently outperforms Supcon across all four tasks, indicating that additional information is learned from arbitrary pairs even with bad pseudo labels.
@@ D.2 Training Detail
config SimLAP SimLAP Hydra-MoCo